OpenStudy (anonymous):

if the area of a triangle is 24 squared inches and the height is 4 inches. how long is the base of the triangle?

OpenStudy (saifoo.khan):

Area = 1/2 * b * h 24 = 1/2 * 4 * b 24 = 2 * b b = 24/2 b = 12in.

OpenStudy (anonymous):

\[Area_{\Delta} = \frac{1}{2}(Base\times Height)\] \[\implies 24 = \frac{1}{2}(Base\times 4)\]